OpenStudy (anonymous):

\[\cos(0.5A) . \cos(0.25A) . \cos(0.125A) ........n = ?\]

OpenStudy (anonymous):

You know the answer?

OpenStudy (anonymous):

Yup... \[\frac{ 2^{-n}sinA }{ Sin \frac{ A }{ 2^n} }\]

OpenStudy (anonymous):

Ohh. Its not till infinity, My bad, Then it's simple. Hint: Write it as: \[[2\sin(A/2^n)\cos(A/2^n)\cos(A/2^{n-1})......\cos(A/2)]/2\sin(A/2^n)\]

OpenStudy (anonymous):

Now use: \[2sinAcosA=\sin2A \]

OpenStudy (anonymous):

And watch the terms combining :)

OpenStudy (anonymous):

Thats Helps